OpenStudy (anonymous):

will be rewarded Fan + Medal write the expression using exponents 3x3x3x3 1.3/3 3. 4/3 4. 3/5

OpenStudy (anonymous):

Is the division sign suppose to be a hat, 3^3 or lets put it in LaTeX \[3^3, ~~~~4^3,~~~~3^5\]

OpenStudy (anonymous):

You're missing one :)

OpenStudy (anonymous):

An option, but anyways do you have any idea?

OpenStudy (anonymous):

Yes, but hey do you have any idea how to do this problem?

OpenStudy (anonymous):

Well for example, if you have something as \[x^2 \times x^5 \implies x^{2+5} \implies x^7\] so are you able to compare this with your problem?

OpenStudy (anonymous):

If you have \[m \times m \times m \implies m^3\] what can you say about your problem?

OpenStudy (anonymous):

It's just a variable, it can represent any number, I was just showing an example.

OpenStudy (anonymous):

Mhm symbol for a number we don't know

OpenStudy (anonymous):

Basically what I'm trying to tell you is, when you multiply numbers/ variables, you add the total amount of numbers/variables for the exponent. So what will be your answer?

OpenStudy (anonymous):

It tells us how many times to use the number when we are multiplying

OpenStudy (anonymous):

Try out this problem, with the examples and definition I've gave you, take your time, and get back to me once you think you know the answer :), it's alright if you're wrong just make sure to make an attempt!

OpenStudy (anonymous):

\[3 \times 3 \times 3 \implies 3^3\] do the same now with your problem...